OpenStudy (anonymous):

\lim_{x\rightarrow 1}\frac{\sqrt[4]{2x-1}+\sqrt[5]{x-2}}{x-1}

OpenStudy (turingtest):

\[\lim_{x\rightarrow 1}\frac{\sqrt[4]{2x-1}+\sqrt[5]{x-2}}{x-1}\]

OpenStudy (anonymous):

yes.this is the question i want to ask

OpenStudy (anonymous):

i thought i answered this one. maybe wasn't clear

OpenStudy (turingtest):

split it up, right sat ?

OpenStudy (turingtest):

oh...

OpenStudy (anonymous):

but i don't understand what u 've said

OpenStudy (anonymous):

lets put \[f(x)=\sqrt[4]{2x-1}+\sqrt[5]{x-2}\]

OpenStudy (anonymous):

then \[f(1)=0\] right?

OpenStudy (anonymous):

and one definition of the derivative is \[f'(a)=\lim_{x\rightarrow a}\frac{f(x)-f(a)}{x-a}\]

OpenStudy (anonymous):

so what you have in this case is \[f'(1)=\lim_{x\rightarrow 1}\frac{\sqrt[4]{2x-1}+\sqrt[5]{x-2}}{x-1}\]

OpenStudy (bahrom7893):

Satellite quick question, what was the answer to yours? P(Adam wins both) - P(Adam loses on first)? That doesn't make sense to me. And the question's bugging me

OpenStudy (anonymous):

so your real job is to find \[f'(x)\] and then replace x by 1

OpenStudy (anonymous):

i mean you still have to do it, that is you have to find \[f'(x)\] but that is not too hard, using the power rule. you get \[f'(x)=\frac{1}{2}(2x-1)^{-\frac{3}{4}}-\frac{1}{5}(x-2)^{-\frac{4}{5}}\]

OpenStudy (anonymous):

so \[f'(1)=\frac{1}{2}(2-1)^{-\frac{3}{4}}-\frac{1}{5}(1-2)^{-\frac{4}{5}}\]

OpenStudy (anonymous):

@lazypig, if this is not clear let me know

OpenStudy (anonymous):

yes.I still don't understand what u've said.

OpenStudy (anonymous):

ok lets go slow

OpenStudy (anonymous):

suppose i have a simpler problem,namly \[f(x)=x^2\]and i want \[f'(3)\]

OpenStudy (anonymous):

one way to solve is say, the derivative of \[f'(x)=2x\] and so \[f'(3)=6\]

OpenStudy (anonymous):

the other is by the definition \[f'(3)=\lim_{x\rightarrow 3}\frac{f(x)-f(3)}{x-3}\] \[f'(3)=\frac{x^2-9}{x-3}\] but since i already know \[f'(3)=6\] i know that that limit must also be 6 of course i can always compute it directly

OpenStudy (anonymous):

now you have \[\lim_{x\rightarrow 1}\frac{\sqrt[4]{2x-1}+\sqrt[5]{x-2}}{x-1}\] so i view this the same way

OpenStudy (anonymous):

i say \[f(x)=\sqrt[4]{2x-1}+\sqrt[5]{x-2}\]

OpenStudy (anonymous):

\[f'(1)=\lim_{x\rightarrow 1}\frac{\sqrt[4]{2x-1}+\sqrt[5]{x-2}-f(1)}{x-1}\]

OpenStudy (anonymous):

but this time i do not want to compute this limit directly because i do not know how. so instead i think lets just take the derivative, replace x by 1 and see what i get, because that is the answer. just like if i have \[\lim_{x\rightarrow 3}\frac{x^2-9}{x-3}\] i can either compute this limit directly (which is easy) or i can think "this is the derivative of x^2 at 3, and that is 6"

OpenStudy (anonymous):

i am not sure that the previous made any sense, but to be brief the gimmick is to recognize this as a derivative, that is the derivative of the numerator, evaluated at x = 1. so take the derivative and then replace x by 1

OpenStudy (anonymous):

I think your answer is quite logical.but i'll solve it by another way.i had just thought of it few minutes ago

OpenStudy (anonymous):

ok. whatever works. you can also use l'hopital, but since the derivative of the denominator is 1, this is identical to finding the derivative of the numerator and replacing it by 1 you cannot do this by multiplying by the conjugate, because the radicals are different

OpenStudy (anonymous):

i find out the answer is 7/10.if you want to know my solution i'll show it

OpenStudy (anonymous):

\[b=\sqrt[5]{x-2}->x=b ^{5}+2\]\[a=\sqrt[4]{2x-1}->x=(a ^{4}+1)\div2\] \[\lim_{x \rightarrow 1}\left(\begin{matrix}\sqrt[4]{2x-1}+\sqrt[5]{x-2} \\ x ^{2}\end{matrix}\right)=\lim_{x \rightarrow 1}\left(\begin{matrix}\sqrt[4]{2x-1}-1 \\ x-1\end{matrix}\right)+\lim_{x \rightarrow 1}\left(\begin{matrix}\sqrt[5]{x-2}+1 \\ x-1\end{matrix}\right)=P+Q\] \[P=\lim_{X \rightarrow 1}\left(\begin{matrix}\sqrt[4]{2x-1}-1 \\x-1 \end{matrix}\right)=\lim_{a \rightarrow 1}\left(\begin{matrix}2(a-1) \\ a ^{4}-1\end{matrix}\right)=\lim_{a \rightarrow 1}\left(\begin{matrix}2 \\ \left( a ^{2}+1 \right)\left( a+1 \right)\end{matrix}\right)=\left(\begin{matrix}1 \\ 2\end{matrix}\right)\] \[Q=\lim_{x \rightarrow 1}\left(\begin{matrix}\sqrt[5]{x-2}+1 \\ x-1\end{matrix}\right)=\lim_{b \rightarrow -1}\left(\begin{matrix}b+1 \\ b ^{5}+1\end{matrix}\right)=\lim_{b \rightarrow -1}\left(\begin{matrix}1 \\ b ^{4}-b ^{3}+b ^{2}-b+1\end{matrix}\right)=\left(\begin{matrix}1 \\ 5\end{matrix}\right)\] \[P+Q=7/10\]